A ball is tossed at a speed of 25 m/2 horizontally off a cliff of a height h. The ball lands 71.4 m away from the base of the cliff. What was the height of the cliff?

Answer options:
a. 14m
b. 28m
c. 40m
d. 80m
e. 120m
1
Expert's answer
2020-12-06T17:27:12-0500

"t=71.4\/25=2.856(s)"


"h=gt^2/2=9.81\cdot 2.856^2/2=40(m)"